### v1

# An Exploratory Case Study of Query Plan Representations

Abstract.
In database systems, a query plan is a series of concrete internal steps to execute a query. Multiple testing approaches utilize query plans for finding bugs. However, query plans are represented in a database-specific manner, so implementing these testing approaches requires a non-trivial effort, hindering their adoption. We envision that a unified query plan representation can facilitate the implementation of these approaches. In this paper, we present an exploratory case study to investigate query plan representations in nine widely-used database systems. Our study shows that query plan representations consist of three conceptual components: operations, properties, and formats, which enable us to design a unified query plan representation. Based on it, existing testing methods can be efficiently adopted, finding 17 previously unknown and unique bugs. Additionally, the unified query plan representation can facilitate other applications. Existing visualization tools can support multiple database systems based on the unified query plan representation with moderate implementation effort, and comparing unified query plans across database systems provides actionable insights to improve their performance. We expect that the unified query plan representation will enable the exploration of additional application scenarios.
Case Study, Testing, Query Plan, Unified Representation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Empirical studiesccs: Software and its engineering Specification languagesccs: Information systems Database utilities and tools
1. Introduction
Database Management Systems (DBMSs) are fundamental software systems used to store, retrieve, and run queries on data. They are used in almost every computing device (Website, 2022b, c, a); thus, any bug has potentially severe consequences. Recently, automated testing approaches for DBMSs have gained broad adoption. Query Plan Guidance (QPG(Ba and Rigger, 2023) guides test case generation towards diverse query plans aiming to expose more bugs, Cardinality Estimation Restriction Testing (CERT(Ba and Rigger, 2024) identifies performance bugs by examining query plans, and Mozi (Liang et al., 2024) checks the consistency across query plans of the same query to find incorrect query optimizations.
While these approaches have been effective and have found more than 200 previously unknown and unique bugs, implementing them requires a non-trivial effort, as they require DBMS-specific logic to parse and process query plans. A query plan refers to a series of concrete steps to execute a query written in a declarative language, such as Structured Query Language (SQL). DBMSs expose query plans in various formats (e.g., in textual format), in which case we refer to them as serialized query plans. Unlike query languages, for which widely-used, standardized (Guagliardo and Libkin, 2018), and formalized (Guagliardo and Libkin, 2017) languages exist, such as SQL (Website, 2023d), the formats in which serialized query plans are exposed are non-standardized and DBMS-specific. For example, a predicate in the WHERE clause of SQL corresponds to a concrete step to filter data in the query plans of TiDB (Bousquet et al., 2001), but corresponds to a property of another step to scan tables in the query plans of PostgreSQL (Guliato et al., 2009). We refer to the different ways in which serialized query plans are represented as query plan representations. Considering that hundreds of DBMSs exist, [^1] https://dbdb.io/ lists 1001 DBMSs as of August 2024. implementing the above testing methods requires significant effort as they need to account for differences in query plan representations, thus significantly hindering the effectiveness of the above approaches.
We envision that a unified query plan representation would remove the roadblock to implementing the above testing approaches. In this work, we systematically study query plan representations. We present an exploratory case study (Runeson et al., 2012), which is a method to investigate a phenomenon in depth, including both qualitative and quantitative research methods. We collected documents, source code, and third-party applications of the query plans in nine popular DBMSs across five different data models, and summarized the commonalities and differences of query plan representations. Our study shows that query plan representations are based on three conceptual components: operations, properties, and formats. Based on the study, we designed a unified query plan representation, which allows the representation of all conceptual components we studied from DBMS-specific query plans, so that we can easily apply automated testing methods to multiple targets.
To demonstrate the utility of the unified query plan representation, we implemented a prototype called UPlan to maintain the representation. Currently, we implemented converters that convert a DBMS-specific serialized query plan to a unified representation; in the long term, we hope that DBMSs will directly expose such a unified representation, rendering such converters unnecessary. We re-implemented QPG and CERT in a DBMS-agnostic way based on UPlan, enabling the large-scale adoption of both testing methods. In our evaluation, we considered three DBMSs: MySQL, PostgreSQL, and TiDB. We found 17 previously unknown and unique bugs, and one bug in the original QPG implementation for parsing TiDB’s query plans. Additionally, we found that UPlan can facilitate other applications based on query plans. First, we modified a visualization tool to support our unified query plan representation. We show that an existing DBMS-specific visualization tool can support five DBMSs through UPlan with only moderate implementation effort. Second, the unified query plan representation enables the comparison of query plans in different DBMSs, which provides actionable insights for query optimization.
Overall, we believe that our work is significant for the software engineering research community, because we focus on how software developers can develop applications based on serialized query plans. We believe that our unified representation reduces the effort to build applications, including testing, on serialized query plans. We include all study results and the prototype in our supplementary materials. We provide a comprehensive website, which provides supportive additional information including illustrative examples, explanations of studied query plans, and example applications. We make the following contributions:
  • A study of query plan representations and results that are available at https://unifiedqueryplan.github.io/, allowing both practitioners and researchers to study the results.
  • A proposal and a reusable library UPlan for a unified query plan representation.
  • Three applications, including testing, of the unified representations on serialized query plans.
2. Background
Query optimization. DBMSs consider a variety of potential execution strategies for a query written in a declarative language (Mella et al., 2019), and each execution strategy includes specific steps to execute the query. Query optimization is the process of determining an execution strategy, and typically includes these steps: parsing queries into logical query plans that represent the logical steps to execute the query, including operations such as join, converting them into physical query plans, which represent the executable operations to execute the query, such as hash join, and determining a physical query plan to execute.
Query plans. Query plans, typically short for physical query plans, are organized as a Directed Acyclic Graph (DAG) (Neumann, 2005), in which a step outputs data to another step as input followed by directed edges. As query optimizations are specific to the storage engines and query execution engines, query plans are non-standardized. By executing a query with a specific prefix, such as EXPLAIN for most relational DBMSs that support SQL, we can obtain a serialized query plan in a DBMS-specific query plan representation. In this paper, we study various query plan representations. We did not study logical query plans as they usually are not exposed by DBMSs.
3. Query Plan Case Study
We adopted an exploratory case study as the method to investigate query plan representations. The case study, as an empirical method, is used for investigating a contemporary phenomenon in depth and within its real-world context (van Rosmalen et al., 2014) (the “case”). An exploratory case study is a specific case study that generates new questions, propositions, or hypotheses during the study. In this paper, we chose this method, because we wanted to gain an in-depth understanding of how query plans are represented within mature DBMSs. We followed common guidelines of case study research (Runeson et al., 2012) to design and conduct this study.
3.1. Case Study Design
Table 1. The studied nine DBMSs ranging from various data models, development modes, and release dates.
DBMS Version Data Model Release Rank
InfluxDB (Anih et al., 2020) 2.7.0 Time-series 2013 28
MongoDB (Ouyang et al., 2022) 6.0.5 Document 2009 5
MySQL (Egea et al., 2010) 8.0.32 Relational 1995 2
Neo4j (Rabuzin et al., 2022) 5.6.0 Graph 2007 21
PostgreSQL (Guliato et al., 2009) 14.7 Relational 1989 4
SQL Server (Website, 1989) 16.0.4015.1 Relational 1989 3
SQLite (Pawlaszczyk, 2022) 3.41.2 Relational 1990 10
SparkSQL (Meng et al., 2016) 3.3.2 Relational 2014 33
TiDB (Bousquet et al., 2001) 6.5.1 Relational 2016 79
Objectives and research questions. The goal of this study was to investigate the phenomenon of non-standardized query plan representations within real-world DBMSs. We aim to achieve this goal by answering the following research questions (RQs):
  • RQ1
    How are serialized query plans represented?
  • RQ2
    Do query plan representations share a common conceptual basis?
Case selection. The case study of this paper is characterized as single-case (Runeson et al., 2012): the query plan representation is the case, while different DBMSs are the units of the analysis. Table 1 shows the DBMSs that we selected for the study. To conduct a representative and comprehensive study, we made a diverse selection of DBMSs. First, we chose DBMSs of various data models: relational, document, graph, and time-series data models, based on the assumption that the DBMSs of different models might have different query plan representations. The relational data model is the most widely-used one (Codd, 1970), and other non-relational models, which are also called NoSQL models, are widely used for maintaining unstructured or semi-structured data (Strauch et al., 2011). Then, we chose both classic and new-generation DBMSs (Pavlo and Aslett, 2016) ranging from release years from the 1980s to the 2010s. The architectures of the chosen DBMSs include standalone DBMSs and an embedded DBMS, SQLite, which runs in the same process as the application. We included both a commercial DBMS, SQL Server, and open-source DBMSs. Apart from conventional DBMSs, we included an analytics engine for large-scale data processing, SparkSQL, which also optimizes queries. To choose widely-used DBMSs for study, we chose them referring to the DBMS ranking website [^2] https://db-engines.com/en/ranking as of August 2024. as shown in the column Rank. We chose the latest release version of each DBMS.
Data collection. We collected data from multiple sources: documents, source code, officially integrated development environments (IDEs), and third-party applications based on query plans. The use of multiple data sources allowed us to perform data source triangulation (Runeson et al., 2012), that is, we could confirm the study results from the above four different types of data sources. We included the detailed lists of the data sources in the supplementary materials for reference.
Data analysis. For each DBMS, we first examined its official documents describing its query plan representation. If the source code was available, we inspected the relevant source code to gain a better understanding of query plan representations. We also ran official test cases in the official IDEs to observe query plan representations in real-world test cases. For the DBMSs that had open-source query plan applications, we further examined how query plan representations are explained and utilized from a third-party perspective. To answer RQ1, we analyzed the query plan representations from the above data sources qualitatively. To answer RQ2, we qualitatively identified conceptual components in the query plan representations, and quantitatively compared these components. To satisfy observation triangulation (Runeson et al., 2012), one author conducted the study, and another author validated the finding against the raw data.
Listing 1: Query plan examples of PostgreSQL and SQLite in text format. Some properties are truncated by … due to page limitation.
1CREATE TABLE t0 (c0 INT);
2CREATE TABLE t1 (c0 INT);
3CREATE TABLE t2 (c0 INT PRIMARY KEY);
4INSERT INTO t0 SELECT * FROM generate_series(1,1000000);
5INSERT INTO t2 SELECT * FROM generate_series(1,100);
6SELECT t1.c0 FROM t0 INNER JOIN t1 ON t0.c0 = t1.c0 WHERE t0.c0 < 100 GROUP BY t1.c0 UNION SELECT c0 FROM t2 WHERE c0 < 10;
7---------------------PostgreSQL-------------------------------------SQLite---------------@HashAggregate@ (cost=62998.82..63009.32 rows=1050...) ‘--@COMPOUND QUERY@ Group Key: t1.c0--LEFT-MOST SUBQUERY
8 ->Append (cost=27150.40..62996.20 rows=1050 width=4) --SCAN t0
9 ->Group (cost=27150.40..62949.08 rows=200 width=4) --SEARCH t1 USING INDEX(c0)...
10 Group Key: t1.c0 --USE TEMP B-TREE FOR GROUP BY
11 ->Gather Set (cost=27150.40..62948.08 rows=400...) --UNION USING TEMP B-TREE
12 Workers Planned: 2 --SEARCH t2 USING INDEX ...
13 ->Group (cost=26150.38..61901.89 rows=200 width=4)
14 Group Key: t1.c0
15 ->Set Join (cost=26150.38..56906.48 rows=1998164 width=4)
16 Set Cond: (t0.c0 = t1.c0)
17 ->Sort (cost=25970.60..26362.39 rows=156719 width=4)
18 Sort Key: t0.c0
19 ->Parallel Seq Scan on t0 (cost=0.00..10301.95 rows=156719 width=4)
20 Filter: (c0 < 100)
21 ->Sort (cost=179.78..186.16 rows=2550 width=4)
22 Sort Key: t1.c0
23 ->Seq Scan on t1 (cost=0.00..35.50 rows=2550 width=4)
24 ->Bitmap Heap Scan on t2 (cost=10.74..31.37 rows=850 width=4)
25 Recheck Cond: (c0 < 10)
26 ->Bitmap Index Scan on t2_pkey (cost=0.00..10.53 rows=850 width=0)
27 Index Cond: (c0 < 10)
28Planning Time: 0.124 ms
3.2. Findings Overview
We found that the studied DBMSs share three conceptual components: operations, properties, and formats. As described in section 2, query plans are considered DAGs. In practice, serialized query plans are usually organized in a tree structure, and each node has a key. The keys typically refer to operations, which are concrete steps executed by DBMSs to retrieve, process, or output data in response to a query, such as Full Table Scan, which refers to the step to scan an entire table. Each operation is associated with zero or multiple properties, which involve operation-related information, such as row, which refers to the estimated number of rows returned. Apart from operation-associated properties, plans also have properties associated with them, such as planning time, which refers to the time to generate the query plan. A special case is the key Filter in the TiDB query plans. It represents the condition that the output data from its child node satisfies, so we deem it as a property instead of an operation. DBMSs typically allow serializing query plans in various formats, such as text, table, JSON, and XML.
1 shows two examples of query plan representations of PostgreSQL and SQLite in a textual format. Lines 1–5 show the SQL statements that create and populate the tables. The function generate_series generates data to populate tables t0 and t2. For the query in line 5, the serialized query plan in PostgreSQL is shown in the left part, and that in SQLite is shown in the right part. The bold texts denote operations, and the non-bold texts denote properties. For example, the operation HashAggregate in line 8 has the properties cost, rows, width, and Group Key. The tree structure is denoted through hierarchical indents, in which the operation with longer indents is a child of the operation with short indents. Although both DBMSs are based on the relational data model, and support a textual format, their query plan representations are significantly different.
Table 2. The number of operations and properties in query plan representations.
Operations Properties
DBMS Producer Combinator Join Folder Projector Executor Consumer Sum Card. Cost Conf. Status Sum
InfluxDB 0 0 0 0 0 0 0 0 5 0 0 1 6
MongoDB 14 9 0 5 3 10 3 44 16 5 18 12 51
MySQL 15 3 2 1 0 2 0 23 3 6 3 10 22
Neo4j 18 11 43 6 3 17 13 111 3 3 12 7 25
PostgreSQL 18 8 3 3 0 9 1 42 8 17 42 40 107
SQL Server 15 3 3 3 0 16 19 59 4 4 7 3 18
SQLite 3 6 3 0 0 5 0 17 0 0 3 0 3
SparkSQL 7 1 2 6 0 43 18 77 11 11 0 0 22
TiDB 19 6 7 5 1 13 5 56 2 5 4 1 12
Avg: 12 5 7 3 1 13 7 48 6 6 10 8 30
3.3. Operations
Identification. We identified operations from the source code of all DBMSs, except for SQL Server, whose source code is not public. MongoDB, MySQL, PostgreSQL, and TiDB specify operations in enumeration variables or lists. Neo4j and SparkSQL define each operation as a class or structure. SQLite defines operations as strings that are passed to the query plan generation process. We found that only SQL Server and Neo4j provided detailed documents for operations, while the other DBMSs’ documents had incomplete lists of operations, and relied on illustrative examples.
Classification. We classified operations into seven categories, as shown in the left part of Table 2. On average, every DBMS defines 48 operations in query plans. Neo4j has the most operations, while InfluxDB has no operations. We noticed that Neo4j requires more operations on nodes and relationships of the graph data model, such as the operation to set node properties (Website, 2024e), and SQLite is a lightweight DBMS with a limited number of operations, such as lacking the operations for creating tables. Overall, NoSQL DBMSs have more operations than relational DBMSs. An exception is InfluxDB, which does not define operations. InfluxDB’s operations are disregarded in query plans due to the limited set of operations supported by the single-entity time-series data.
Terminology. We define the operation categories through the following terms. Recall that data models organize data in entities and attributes, so we represent an entity $E=\{a_{1},a_{2},...,a_{n}\}$, in which $a$ is an attribute, and represent a multiset [^3] A collection of elements in which elements may occur more than once. of entities $D=\{E_{1},E_{2},...,E_{n}\}$. $O_{c}(input)=output$ represents that an operation $O$ of the category $c$ receives input from children operations and produces output to the parent operation. Considering that many operations might apply to varying numbers of children, we chose to model the input from all children as a single input. We give a detailed explanation of each operation category as follows.
Producer. The Producer category consists of the operations that retrieve data from storage or return constants instead of from children’s operations. We denote them as $O_{Producer}()=D_{o}$. The operations in the Producer category are data sources of queries, so they are typically leaf nodes of query plans. For example, in 1, the operation SEARCH in the right part of line 11 represents a full table scan, and Bitmap Heap Scan in line 26 represents a data scan from bitmaps in heap memory. Six of nine DBMSs define more than ten operations in the Producer category, because reading data is usually expensive, and thus reads are customized for different scenarios aiming to improve efficiency. For example, indexes (Guttman, 1984; Lehman and Carey, 1986) can be used to efficiently read data.
Combinator. The Combinator category consists of the operations that change the permutation and combination of entities, such as sort and union, with no changes to attributes. We denote them as $O_{Combinator}(\{E_{i}|E_{i}\in D_{i}\})=\{E_{o}|E_{o}\in D_{o},D_{o}\subseteq D_{i}\}$. In 1, the operation Append of PostgreSQL merge entities from different children operations to a single set by the operations in lines 11 and 26, and is associated with query clause UNION in line 6. To execute a similar functionality in SQLite, the operations COMPOUND QUERY and UNION combine data objects by the operations in the right part of lines 8 and 13, and both operations also belong to Combinator category.
Join. The Join category consists of the operations that generate new entities by recombining attributes. We denote them as $O_{Join}(\{E_{i}|E_{i}\in D_{i}\})=\{E_{o}|\forall a_{o}\in E_{o},a_{o}\in D_{o},D_{o}\subseteq D_{i}\}$. In 1, the operation Set Join in line 17 combines two ordered data from the operations in lines 19 and 23 based on the common fields t0.c0 and t1.c0. MongoDB has no Join operations, because it includes only a single document entity for querying and lacks support for combining data from multiple documents. The higher number of Join operations in Neo4j is because we classified the operations on the edges of the graph data model as belonging to the Join category. In the graph data model, edges establish relationships between nodes, and a broader range of operations can be performed on the edges. For example, in Neo4j, executing the simple query MATCH ()-[r]->() WHERE r.title ENDS WITH ’developer’ RETURN r retrieves the relationships whose properties satisfy r.title ENDS WITH ’developer’. The corresponding query plan is shown in Figure 1. Each line in the table represents an operation and associated properties, and the content outside the table is plan-associated properties. The query plan specifies scanning the relationships, which indicates both nodes (i.e., entities) each relationship connects, so the operation UndirectedRelationshipIndexContainsScan belongs to Join.
Figure 1. An example query plan of the Neo4j operations of the Join category.
Planner COST
Runtime version 5.10
Operator Rows
+ProduceResults 8
+UndirectedRelationshipIndexContainsScan 8
Total database accesses: 5, total allocated memory: 184
Folder. The Folder category consists of the operations that derive new entities from a set of entities. Formally, we denote them as $O_{Folder}(\{E_{i}|E_{i}\in D_{i}\})=\{E_{o}|E_{o}=\psi(D_{o}),D_{o}\subseteq D_{i}\}$, in which $\psi()$ denotes a function, such as $MAX()$. In 1, the operations HashAggregate and Group are in the Folder category and represent data aggregation and grouping, respectively. SQLite does not define operations in the Folder category, but shows similar information in properties together with the operations in the Producer category. The operations in the Folder denote DBMSs’ data transformation capability, so most DBMSs support operations in the Folder category.
Projector. The Projector category consists of the operations that remove attributes from all entities. Formally, we denote them as $O_{Projector}(\{E_{i}|E_{i}\in D_{i}\})=\{E_{o}|E_{o}\subseteq E_{i}\}$. No operation in 1 belongs to this category, and 6 of 9 DBMSs have no operations in this category. We observed that these operations correspond to SELECT clauses, and they are not explicitly denoted in query plans.
Executor. The Executor category consists of the operations that make no change to entities and attributes. We denote them as $O_{Executor}({D_{i}})=\{D_{o}|D_{o}=D_{i}\}$. In 1, the operation Gather Set in line 15 merges the data from the operation Group running in other computing nodes for a distributed architecture. DBMSs define these operations to cater to various designs and goals. For example, PostgreSQL defines the operation MEMORIZE to cache the output from node children into memory to speed up processing. We classified these operations into the Executor category. The average number of operations in the Executor category is 13, and SparkSQL has significantly more operations, 43, in the Executor category than others, because it defines multiple operations to interact with other components, such as the Python library pandas.
Consumer. The Consumer category consists of operations that have no output. We denote them as $O_{Consumer}(D_{i})=\{\}$. Apart from queries, which, in the SQL, are SELECT statements, DBMSs also support other statements, such as CREATE and UPDATE in SQL. DBMSs also expose query plans for these statements, and name them as execution plans for wider usage (Website, 2024b). The operations of the Consumer category usually modify stored data or system variables. For example, SparkSQL uses the operation SetCatalogAndNamespace to control a particular system variable, and we assign these operations to the Consumer category.
3.4. Properties
Identification. Each property is associated with either an operation or a query plan, and the available properties are statically encoded as strings near the generation processes of associated operations or query plans in the source code. InfluxDB’s query plan representation includes only a list of plan-associated properties, while other DBMSs include both plan-associated and operation-associated properties. In 1, PostgreSQL’s operations have various general properties enclosed in brackets, along with operation-specific properties in the subsequent lines. At the bottom of the serialized query plan, the property Planning Time is plan-associated and represents the time to produce the query plan. In the documentation, similar to operations, only SparkSQL provides a comprehensive list of properties, while other DBMSs only show examples of properties. We explain that it is difficult to maintain the documentation of properties, which are diverse and evolving over versions. To maintain the information of properties, some third-party tools, like pgMustard, maintain a curated list of properties with accompanying explanations, but are usually commercial.
Classification. We identified four categories of properties, as shown in the right part of Table 2. On average, every DBMS defines 30 properties. PostgreSQL has the most properties, since it includes many fine-grained properties. For example, it defines three properties to show the status of parallel computations: worker number, worker launched, and worker planned, while other DBMSs provide at most one property for parallel computation.
Cardinality. The Cardinality category consists of the numeric properties that denote the estimated data size returned by operations. The properties in this category can be associated with operations of any category or the serialized query plan as a whole. In 1, the properties rows and width belong to the Cardinality category and represent the estimated number of returned rows and width. These estimates are derived from statistical information (Ioannidis, 2003) that DBMSs collect, such as the total number of rows and maximum values. Query plans with lower estimated cardinalities are more likely to be selected for execution during cost-based query optimization. Some DBMSs, such as MySQL, provide more fine-grained information about the number of rows that are read and returned. As a lightweight DBMS, SQLite uses simple heuristics to estimate cardinalities, and omits properties in the Cardinality category.
Cost. The Cost category consists of the numeric properties that denote the estimated resource consumption. The properties in this category can be associated with operations of any category or the serialized query plan as a whole. In 1, the property cost is in the Cost category, and the two numbers of cost denote the cost scores of starting and finishing the associated operation by estimating the total consumption of disk and CPU. As in the Cardinality category, SQLite lacks properties in the Cost category.
Configuration. The Configuration category consists of the properties that configure the operations’ parameters, and their values are configuration options which are usually strings or boolean values. The properties in the Configuration category can be associated with operations in any category or the serialized query plan, and are typically specific to operations. In 1, PostgreSQL’s properties Group Key, Set Cond, Sort Key, Recheck Cond, Index Cond, Filter are in the Configuration category and are specific to the associated operations to show the keys used to group, the condition to join, the key to sort, the condition to check, the index condition, and the predicate to exclude data, respectively. SQLite’s property USING INDEX denotes the index condition.
Status. The Status category consists of the properties of run-time status, and their values are runtime metrics which are usually strings or numbers. These properties can be associated with operations in any category or the serialized query plans, and typically differ depending on the operations they are attached to. In 1, the property Workers Planned is in the Status category and shows the number of available computing nodes to execute the associated operation. As another example, TiDB defines the property taskType to show the name of the computing nodes that the operation is assigned to execute. The properties in the Status category show running status, and are determined by the execution environment, while the properties in the Parameters are usually decided by queries. The properties in both Status and Configuration categories are customized, and thus are typically different across DBMSs, while the properties in other categories share similar semantics or functionalities across DBMSs.
3.5. Formats
Table 3. The officially supported formats of query plans.
Natural Structured
DBMS Graph Text Table JSON XML YAML
InfluxDB
MongoDB
MySQL
Neo4j
PostgreSQL
SQL Server
SQLite
SparkSQL
TiDB
DBMSs serialize query plans to various formats for different purposes. The formats are typically controlled by a specific configuration in queries, such as for PostgreSQL, the statement EXPLAIN (FORMAT JSON) SELECT…serializes the query plan representation to JSON format. We classified all formats into two categories: natural formats, which are optimized for readability, and structured formats, which are optimized for machine reading. We also consider the graph formats that are supported in official IDEs.
Table 3 shows the different formats of query plan representations. Overall, DBMSs support more formats in the natural category rather than the structured category, suggesting that DBMSs prioritize readability over machine processing. Due to the lack of a standard, none of the formats is supported by all DBMSs. For the same DBMS, the formats in the natural category usually include less information than the formats in the structured category. For example, in 1, the property Parent Relationship represents how the associated operation passes data to another operation. This property is ignored in the text format of the natural category, but is shown in the JSON format of the structured category. We provide more details of each format as follows.
Natural category. The natural category includes graph, text, and table formats. Query plans are usually serialized as graphs for DBMSs’ IDEs, such as Workbench [^4] https://dev.mysql.com/doc/workbench/en/wb-performance-explain.html for MySQL, Compass [^5] https://www.mongodb.com/docs/compass/current/query-plan/ for MongoDB. Text formats represent query plans as plain text, such as shown in 1. Table formats encode each operation and associated properties in a line, and use line numbers to represent the tree structure of query plans. Graph formats are intuitive to understand, so graph formats are supported by most DBMSs.
Structured category. The structured category includes the JSON, XML, and YAML formats. These formats are standardized and widely used for exchanging data (International, 2013; Bosak, 1997). JSON is more widely supported by DBMSs than other structured formats, and PostgreSQL supports all structured formats. Structured formats are not supported by some DBMSs, such as InfluxDB, SQLite, and SparkSQL. SQLite can output a structured format of bytecode, which consists of low-level instructions, not operations and properties, so we do not consider it as a structured query plan representation.
Table 4. Third-party visualization tools for query plans.
Tool DBMSs License
Postgres Explain Visualizer 2 (Website, 2023g) PostgreSQL Open-source
pgmustard (Website, 2023f) PostgreSQL Commercial
pganalyze (Website, 2023e) PostgreSQL Commercial
ApexSQL (Website, 2023a) SQL Server Commercial
Plan Explorer (Orthey et al., 2020) SQL Server Commercial
Azure Data Studio (Website, 2023b) SQL Server Commercial
Dbvisualizer (Website, 2023c) MySQL, PostgreSQL, SQL Server Commercial
Visualization. Apart from the graph formats in official IDEs, third-party visualization tools show query plans based on structured formats to enhance the readability of query plans. Table 4 shows the visualization tools we found for the studied DBMSs. Six of the seven tools are commercial, suggesting the value of understanding query plan representations for developers. Building these tools requires non-trivial effort, because a tool is specific to a DBMS.
4. Unified Query Plan Representation
Our study in section 3 shows that query plan representations share the same conceptual basis, which is why we propose a unified query plan representation that is:
  1. (1)
    complete, to include all information of a query plan,
  2. (2)
    general, to support various DBMSs we studied,
  3. (3)
    extensible, to support DBMSs we did not study.
Listing 2: The unified query plan representation in EBNF.
1plan ::= ( tree )? properties
2tree ::= node ( ’--children-->’ ’{’ tree (’,’ tree)* ’}’ )?
3node ::= operation properties
4operation ::= ’Operation’ ’:’ operation_category ’->’ operation_identifier
5properties ::= ( property ( ’,’ property )* )?
6property ::= property_category ’->’ property_identifier ’:’ value
7operation_category ::= ’Producer’ ’Combinator’ ’Join’ ’Folder’ ’Executor’ ’Projector’ ’Consumer’
8property_category ::= ’Cardinality’ ’Cost’ ’Configuration’ ’Status’ operation_identifier ::= keyword property_identifier ::= keyword keyword ::= letter ( letter digit ’_’ )* value ::= string number boolean ’null’
9string ::= ’"’ ( letter digit )* ’"’ number ::= ’-’? digit+ boolean ::= ’true’ ’false’
10letter ::= [a-zA-Z]
11digit ::= [0-9]
Figure 2. Railroad diagrams of the symbols tree and node.
4.1. Design
To define and illustrate the unified query plan representation, we adopted Extend Backus Naur Form (EBNF) (Pattis, 1994), which is a metasyntax notation to express context-free grammars. 2 shows the unified query plan representation in EBNF. Following our study in section 3, the unified representation includes the three identified conceptual components of different categories. We define plan as a tree that can have plan-associated properties. Within the tree, a node is defined as an operation and zero or multiple operation-associated properties. operations and properties are key-value pairs including corresponding categories, identifiers, and values. In Figure 2, we also visualized the EBNF into a railroad diagram for an intuitive explanation. To make it concise, we only show the symbols tree and node, which express the high-level structure.
We use a unified naming convention to denote operations and properties in the unified query plan representation. A unified naming convention increases the readability and consistency of the representation while avoiding name collisions. section 3 shows that various operations and properties share similar semantics, so we mapped DBMS-specific names of operations and properties to unified names. For example, we mapped the operation name Seq Scan in PostgreSQL, Table Scan in SQL Server, and TableFullScan in TiDB to Full Table Scan.
4.2. Analysis
We qualitatively analyze whether this design achieves three goals.
Completeness. The unified query plan representation includes the three conceptual components: operations, properties, and formats, which we identified in section 3. Operations and properties are included in the tree of the unified representation, and the unified representation can be serialized into other standard formats, such as JSON and XML, which are used in query plan representations.
Generality. The unified query plan representation supports the query plan representations of the nine DBMSs that we studied. InfluxDB’s query plan includes a list of properties without operations, which can be represented by plan-associated properties in the unified representation. For the other DBMSs, we identified operations and properties, mapped them into unified names, and organized them into our unified representation.
Extensibility. The definitions of operations, properties, and categories in the unified query plan representation can be extended or shrunk while keeping forward and backward compatibility. Forward compatibility refers to allowing a system accepts input intended for a later version of itself, and backward compatibility refers to allowing a system accepts input intended for an older version of itself (Wallace, 2007). Both evaluate whether the applications based on the unified query plan representation still work if we update the representation to support more or different DBMSs. To keep forward compatibility, we can add more categories by expanding operation_category to include more category names, and add more operations and properties whose names comply with the definition of the keyword at line 11 at 2. An existing application still can parse the revised representation by ignoring the newly added categories, operations, and properties, or handle them in a generic way (e.g., a visualization tool could represent unknown operations using a generic visual shape). For backward compatibility, similarly, an existing application still can parse the old version of the representation, whose categories, operations, and properties are included in the new version of the representation.
5. Applications
We implemented the prototype of a reusable library, UPlan, to maintain the unified query plan representation. While our main motivation is facilitating the implementation of automated testing approaches, the unified representation enables also other applications of visualization and benchmarking:
A.1 Testing.:
The DBMS testing methods QPG and CERT were implemented in a DBMS-specific way due to DBMS-specific query plan representations, and we show how UPlan allows both methods to be implemented in a DBMS-agnostic way.
A.2 Visualization.:
Visualization tools visually display serialized query plans to ease understanding, but are typically specific to a particular DBMS. We show a general visualization tool based on UPlan.
A.3 Benchmarking.:
Benchmarking is an important method to evaluate the performance of DBMSs. We show a case analysis of a comparison of query plan representations using UPlan. We hope that allowing developers to easily compare different DBMSs’ query plans enables them to improve their DBMSs’ query optimization capabilities.
DBMSs. For A.2 and A.3, we used MongoDB, MySQL, Neo4j, PostgreSQL, and TiDB, because they support the JSON format of query plans. JSON is the most widely supported structural format, and it typically includes more detailed information than other formats. For example, JSON may include read and optimization costs, but the text format often only includes overall cost, as it is typically optimized for human readability. Within these DBMSs, we used MySQL, PostgreSQL, and TiDB for A.1 as they are supported by SQLancer. We used the same versions of DBMSs that we studied in the Table 1.
Data set. For A.1, we used SQLancer to generate test cases. For A.2 and A.3, we collected the serialized query plans of the queries from the TPC-H benchmark suite (Website, 2023h) on the five DBMSs. TPC-H comprises 8 tables and 22 queries for relational DBMSs: MySQL, PostgreSQL, and TiDB in our experiments. MongoDB and Neo4j require manual effort to adapt TPC-H as it was not designed for non-SQL languages and non-relational data models. MongoDB is based on a document data model, which lacks support for join operations, so we embedded all entities in one document and rewrote queries 1, 3, and 4 in the MongoDB Query Language (MQL), following a tutorial (Website, 2024c). For Neo4j, which is based on a graph data model, we mapped nodes into rows of the relational data model and edges into foreign keys of the relational data model following another tutorial (Website, 2024d), and rewrote queries 1–14, 16–19 using the Cypher Query Language (CQL) following an example (Website, 2024a).
Implementation. UPlan is a reusable library that consists of around 300 lines of Python code to implement a reusable library, which allows adding or updating operations and properties. The prototype supports serializing query plan representations into JSON as well as text formats, and provides an interface for supporting more formats. We also implemented five customized converters to parse the query plan representations from existing query plan representations to the unified query plan representation, and each parser has around 200 lines of code.
A.1 Testing
We show how UPlan allows implementing the testing methods QPG (Ba and Rigger, 2023) and CERT (Ba and Rigger, 2024) in a DBMS-agnostic way. QPG is a test case generation approach that is guided by query plans; specifically, it mutates a database if no new query plans have been observed for a specific number of randomly generated queries, aiming to subsequently exercise new query plans, and thus exploring “interesting” behaviors. In terms of implementation, evaluating whether a query plan is structurally different from another requires ignoring unstable information, such as random identifiers and the estimated cost in query plans. CERT is a test oracle for finding performance issues by comparing estimated cardinalities, which have to be extracted from query plans. Both methods were implemented in SQLancer, which is a popular and widely used tool for automatically testing DBMSs. The original implementations used DBMS-specific and error-prone methods, such as string matching and substitution, which had to be implemented for every DBMS that was supported. Both approaches support relational DBMSs, but were not applied to the popular open-source DBMSs MySQL and PostgreSQL, as additional DBMS-specific parsers would have been required. Based on UPlan, we implemented general parsers for both methods to support all UPlan-compatible DBMSs. We did not consider Mozi, because its source code is not public.
To evaluate UPlan on QPG and CERT, we applied the general parser to MySQL, PostgreSQL, and TiDB. We ran our revised versions of QPG and CERT for 24 hours and found 17 unique and previously unknown bugs, as shown in Table 5. Developers confirmed 16 of 17 bugs and fixed two bugs. The one bug in PostgreSQL was still waiting for the response from developers, so we did not report more bug reports to avoid burdening developers. 11 of 17 bugs are Critical, Serious, or Major, which represent that the bugs can significantly affect the systems. Additionally, CERT found hundreds of potential bug-inducing test cases in 24 hours, but it is challenging to distinguish their uniqueness, which requires developers’ expertise. To avoid burdening developers, we will report them after the above 17 bugs have been fixed. 3 shows a bug in MySQL found by the test case generated by QPG with UPlan. Note that we identified this bug by the test oracle Ternary Logic Partitioning (TLP) (Rigger and Su, 2020); however, for presentation, we simplified the bug-inducing test case by demonstrating that the same query returns different results in lines 4 and 6 depending on whether the index exists. The cause of the bug was an incorrect table look-up due to the index inserted by the SQL statement in line 5. Using UPlan, we were able to apply QPG to MySQL easily, which was previously incompatible and untested by QPG. This enabled us to find this bug in MySQL.
Table 5. Previously unknown and unique bugs found by QPG with UPlan.
DBMS Found by Bug ID Status Severity
MySQL QPG 113302 Confirmed Critical
MySQL QPG 113304 Confirmed Critical
MySQL QPG 113317 Confirmed Critical
MySQL QPG 114204 Confirmed Serious
MySQL QPG 114217 Confirmed Serious
MySQL QPG 114218 Confirmed Serious
MySQL CERT 114237 Confirmed Performance
PostgreSQL CERT Email Pending Performance
TiDB QPG 49107 Fixed Major
TiDB QPG 49108 Confirmed Major
TiDB QPG 49109 Fixed Major
TiDB QPG 49110 Confirmed Major
TiDB QPG 49131 Confirmed Major
TiDB QPG 51490 Confirmed Moderate
TiDB QPG 51523 Confirmed Moderate
TiDB CERT 51524 Confirmed Minor
TiDB CERT 51525 Confirmed Minor
Listing 3: A bug found by QPG with UPlan.
1CREATE TABLE t0(c0 INT, c1 INT);
2INSERT INTO t0(c1, c0) VALUES(0, 1);
3
4SELECT * FROM t0 WHERE t0.c1 IN (GREATEST(0.1, 0.2)); -- empty result [Uncaptioned image]
5CREATE INDEX i0 ON t0(c1);
6SELECT * FROM t0 WHERE t0.c1 IN (GREATEST(0.1, 0.2)); -- {10} [Uncaptioned image]
We also found that UPlan reduces the risk of implementation bugs for DBMS-specific query plan parsers. We identified an implementation bug in QPG. Specifically, the query plan parser for TiDB failed to exclude random identifiers due to an incorrect parameter for EXPLAIN. With the single implementation for a parser and the unified query plan representation, we have a lower risk of introducing these implementation bugs.
UPlan enables large-scale adoption for testing methods QPG and CERT in a DBMS-agnostic implementation way.
A.2 Visualization
We implemented a visualization tool for serialized query plans by modifying PEV2 (Website, 2023g), a customized query plan visualization tool for PostgreSQL, to use the unified query plan representation. We modified its parser to support identifying the unified query plan representation, and updated its definitions of visualized elements, such as operation names.
Figure 3. Visualized unified query plans of query 1 from TPC-H benchmark.
Using UPlan, our customized PEV2 supports query plan visualization of the five DBMSs in the same implementation by modifying around 800 lines of code. Suppose implementing a new visualization tool for another DBMS also requires at least an additional 800 lines of code, then it would require $800*(5-1)=3,200$ lines of code to support the five DBMSs based on PEV2 without the unified query plan representation. UPlan reduces the effort of building a new visualization tool for serialized query plans and unifies the query plan visualizations across DBMSs.
[…]
Existing DBMS-specific visualization tools could support more DBMSs if they supported our unified query plan representation.
A.3 Benchmarking
In this application, we show how to find potential optimization opportunities by comparing the serialized query plans across DBMSs using the unified query plan representation. Query optimization is a critical process for DBMSs’ performance. To evaluate how effective a query optimization is, existing methods depend on measuring DBMSs’ execution time on standard datasets, such as TPC-H (Website, 2023h) and the Join Order Benchmark (JOB) (Leis et al., 2015). The execution time shows the overall performance difference between various DBMSs, but cannot provide possible reasons for performance gaps across DBMSs. UPlan enables comparing various serialized query plans across DBMSs. Specifically, we collected metrics on the number of operations in DBMSs’ query plan representations and show an example of analyzing the difference.
Table 6. The average number of operations in query plans of the queries from the TPC-H benchmark suite.
DBMS Producer Combinator Join Folder Projector Executor Sum
MongoDB 1.00 0.00 0.00 0.00 1.00 0.00 2.00
MySQL 4.55 0.82 2.77 0.86 0.00 0.27 9.27
Neo4j 0.39 0.78 2.89 0.06 0.72 3.06 7.89
PostgreSQL 3.95 1.32 2.64 1.73 0.00 2.45 12.09
TiDB 4.18 0.82 2.73 1.41 1.77 3.73 14.64
Table 6 shows the average number of operations in each category for the query plans of queries from the TPC-H benchmark suite. We omitted the Consumer category as we did not encounter any such operations. The relational DBMSs, MySQL, PostgreSQL, and TiDB, have more operations than the non-relational DBMSs, MongoDB and Neo4j. It is because relational DBMSs have more operations in the Producer category. To further explain the difference in the Producer category, we looked into query plans and found that each table in relational DBMSs requires at least one operation to read data, while non-relational DBMSs usually read all data in one or two operations. Apart from MongoDB, the other four DBMSs have a similar number of Join operations. The results also show that the query plans of the queries in the TPC-H benchmark suite usually do not cover all categories of operations due to the limited set of queries. The operations of the Producer category are typically expensive for performance, and hence, developers typically aim to analyze and reduce the occurrence of the operations in this category. We explain an example to analyze them as follows.
Figure 4. Variance of the number of Producer operations for each query in TPC-H benchmark across five DBMSs.
Figure 4 shows the variance of the number of operations in the Producer category for each query plan. Among 22 queries, the variances of six queries are more than 5, indicating a significant difference. Queries 2, 8, 5, 7, and 9 have a significant variance due to different data models. For example, for query 2, the DBMSs of relational data models, MySQL, TiDB, and PostgreSQL, have 10, 12, and 9 operations, while the DBMS based on the graph data model, Neo4j, has only 1 operation in the Producer category. Query 11 has a significant variance due to another potential optimization issue, and we explain it as follows.
Listing 4: The unified serialized query plan representations in the text format for query 11 in TPC-H. Underlines represent table names. Predicates in the query and properties in the query plan representations are ignored for brevity.
1SELECT ... FROM PARTSUPP, SUPPLIER, NATION
2WHERE ... GROUP BY ... HAVING ... > (
3 SELECT ... FROM PARTSUPP, SUPPLIER, NATION
4 WHERE ...
5) ORDER BY ...;
6-----------------PostgreSQL-------------------------------------TiDB--------------------
7Combinator->Sort Projector->Project
8 Folder->Aggregate Combinator->Sort
9 Join->Hash Join Folder->Aggregate Hash
10 Producer->Full Table Projector->Project
11 name object: partsupp Join->Index Hash
12 Executor->Hash Row Join->Index Hash
13 Join->Hash Executor->Collect
14 Producer->Full Table Producer->Full Table
15 name object: supplier name object: nation
16 Executor->Hash Row Executor->Collect Order
17 Producer->Full Table Producer->Index-only...
18 name object: nation name object: supplier
19 Folder->Aggregate Executor->Collect Order
20 Join->Hash Join Producer->Index-only...
21 Producer->Full Table name object: partsupp
22 name object: partsupp Producer->Id Scan
23 Executor->Hash Row name object: partsupp
24 Join->Hash Join
25 Producer->Full Table
26 name object: supplier
27 Executor->Hash Row
28 Producer->Full Table
29 name object: nation
4 shows query 11 from the TPC-H benchmark suite and the corresponding serialized query plans of PostgreSQL and TiDB using the unified query plan representations in text format. The query references the three tables PARTSUPP, SUPPLIER, NATION twice in the FROM and HAVING clauses respectively. PostgreSQL uses six table scans, one for each table reference in the original query, while TiDB could optimize the query to use only three scans. The table partsupp is scanned twice in lines 21 and 23, because TiDB reduces the data size of table scanning by retrieving a secondary index before the scan operation. The first scan retrieves indexes only to obtain the row id, which is used for the second scan. Suppose both DBMSs execute the operations in order and the same operation has the same performance overhead in both DBMSs, then the query plan with three table scans is more efficient than the query plan with six table scans. UPlan enables this analysis and provides actionable insight for DBMS developers to improve query performance by reducing three repeated table scans. We reported this issue to the PostgreSQL developers, who confirmed that this case indicates an unsupported optimization by PostgreSQL. One developer considered how this could be supported—”[…] maybe there is some easy way to hook this into the same code used by GROUPING SETS […]”.
Comparing query plans between DBMSs, using the unified query plan representation, provides insights for improving query optimizers.
6. Discussion
Paths to adoption. Developers can make use of the unified query plan representation by customized converters to convert original serialized query plans into the unified query plan representation. For all three applications we showed, we implemented five customized converters, each of which has around 200 lines of code, within one week only. If the converters can be implemented by DBMS developers or experts, who are well-versed in the query plans, it is plausible that a higher-quality converter could be developed in a shorter time. We hope that DBMSs will directly expose query plans in the unified representation in the long term, thus avoiding a conversion.
Additional use cases. We envision several additional use cases that are enabled by our unified query plan representations. Toward a comprehensive evaluation of query optimization, additional metrics could be explored, such as similarity on tree structures (Yang et al., 2005), to compare different DBMSs’ query plans using our unified query plan representation. To find issues in query optimization, we can apply differential testing (McKeeman, 1998) or other methods to compare the unified query plan representations among DBMSs. Query optimization approaches based on machine learning have been proposed that take query plans as input and output suggestions for indexes (Ding et al., 2019), views (Yuan et al., 2020), and join orders (Marcus and Papaemmanouil, 2018; Yu et al., 2020), so our unified query plan representation would allow exchanging training data in different DBMSs to improve the performance of models.
Threats to validity. Our study faces several threats to validity, which denotes the trustworthiness (Neuman, 1991; Runeson et al., 2012) of the results, and to what extent they are unbiased. A major concern is the degree to which the data and analysis depend on the specific researchers. We followed the best practice of triangulation (Runeson et al., 2012), which refers to taking multiple perspectives toward the same object, to increase reliability. For data triangulation, we collected data from multiple sources: documentation, source code, and third-party applications. For observer triangulation, one author conducted the study, and another author validated the findings against the raw data. For methodological triangulation, we used a qualitative method to analyze query plan representations, and a quantitative method to examine and classify the three conceptual components of query plan representations. Furthermore, we have made the process of data collection and analysis publicly available, along with comprehensive study results presented in the supplementary materials of this paper. Another concern is the degree to which our results can be generalized to and across the query plan representations of other DBMSs. We selected representative DBMSs of various data models: relational, document, graph, and time series. The last concern is the degree to which the study really assesses the research questions we aim at. We collected the data ourselves, and our analysis of the semantics may be inconsistent with the intentions of the developers that implemented the query plans.
7. Related Work
Query plan. The most related work is Substrait (substrait io, 2021), which proposes a well-defined, cross-language specification for data-compute operations on structural data. It was designed for interoperability across different languages and systems. One of its use cases is unifying logical query plan representation for DBMSs. It enables executing the query for a specific DBMS across multiple DBMSs by converting queries into unified logical query plans, which are converted to DBMS-specific physical query plans. However, the unified logical query plans, provided by Substrait, cannot facilitate building general applications based on physical query plans such as demonstrated in this work. Logical query plans represent logical steps to execute a given query, and do not include DBMS-specific internal execution steps, which are required for these applications. In this work, we studied how physical query plans are represented in nine real-world DBMSs and proposed a unified representation for physical query plans.
Improving query optimization. Query optimization is a long-standing problem and a core database research topic with a significant body of related work. We discuss several representative works here. Ioannidis et al(Ioannidis, 2003) proposed to sample data in single columns for estimating the cardinality of a query plan. Ilyas et al(Ilyas et al., 2004) proposed to detect correlations across columns, which enables a multi-column sampling (Poosala and Ioannidis, 1997) for a more accurate estimation of cardinality. Lakshmi et al(Lakshmi and Zhou, 1998) proposed using machine learning to estimate cardinality. Estimated cardinality, as well as other estimated costs, are considered in cost models. Babcock et al(Babcock and Chaudhuri, 2005) proposed cost distribution to choose a query plan with the lowest cost in a robust way. Apache published a unified framework Calcite (Begoli et al., 2018) to replace existing query optimizers. Compared to them, we studied various query plan representations, instead of improving query optimization or replacing existing query optimizations.
Benchmark and empirical study of query optimization. Various papers studied query optimizations to understand their performance through empirical studies. Leis et al(Leis et al., 2015) experimentally studied the contributions of query optimization components to overall performance. Ortiz et al(Ortiz et al., 2019) provided an empirical analysis on the accuracy, space, and time trade-off across several machine learning algorithms for cardinality estimation specifically. Harmouch et al(Harmouch and Naumann, 2017) conducted an experimental survey on cardinality estimation, focusing on estimating the number of distinct values. While we also present an empirical study in the context of DBMSs, we focus on query plan representations, rather than the performance, algorithms, and internal components of query optimization.
Applications based on serialized query plans. Several applications based on serialized query plans exist. QE3D (Scheibli et al., 2015) visualizes distributed serialized query plans for an intuitive understanding and analysis. Machine learning algorithms have utilized serialized query plans for query optimization. Yuan et al(Yuan et al., 2020) used machine learning to select optimal views. Yu et al(Yu et al., 2020) used reinforcement learning to determine the join order. Marcus et al(Marcus et al., 2022) and Ryan et al(Marcus et al., 2019) applied machine learning algorithms to generate query plans. Zhao et al(Zhao et al., 2022) used machine learning to convert serialized query plan representations to vector representations to facilitate other machine learning algorithms. In this work, we propose a unified query plan representation, which reduces the effort to build the applications based on serialized query plans.
Standardization. Multiple works were proposed to standardize DBMSs. The most related work is Substrait (substrait io, 2021), which proposes a unified representation for DBMS-agnostic relational algebra, while we studied DBMS-specific query plans. Mitschang (Mitschang, 1988) proposed a unified view of design data and knowledge representation when supporting database systems to non-standard applications, such as Computer-Aided Design (CAD). Ginsburg et al(Ginsburg and Wang, 1992) proposed a unified approach to query sequenced data. Gueidi et al(Gueidi et al., 2021) proposed a unified modeling method for Non-relational DBMSs (NoSQL) to facilitate the applications on NoSQL. Compared to these works, we propose a unified framework for the query plan representations in DBMSs.
8. Conclusion
We have presented an exploratory case study to investigate how query plan representations are in nine widely-used DBMSs. Our study has shown that query plan representations share conceptual components among different DBMSs: operations, properties, and formats. Based on the study, we designed the unified query plan representation to reduce the effort to build applications based on query plans. We implemented a reusable library UPlan to maintain the unified representation, and evaluated it on five DBMSs. The results show that existing testing methods can be efficiently adopted, finding 17 previously unknown and unique bugs. Additionally, existing DBMS-specific visualization tools could support at least five DBMSs by using UPlan with only moderate implementation effort, and UPlan enables comparing query plans in different DBMSs, which provides actionable insights. This paper provides a comprehensive study of query plan representations, and can be used as a reference for other research on serialized query plans. We believe the unified query plan representation provides more opportunities to research serialized query plans in the future.